OpenStudy (anonymous):

PLS FRNDS HELP ME TO SOLVE THIS PUZZLE When can you add 2 to 11 and get 1 as the correct answer?

OpenStudy (anonymous):

try positives and negitive numbers they cancel each other out

OpenStudy (anonymous):

bt how ? i cant able to understand about dis puzzle:P

OpenStudy (anonymous):

is that the exact question?

OpenStudy (anonymous):

so basically 11x +2 =1 solve for x

OpenStudy (anonymous):

yeah its exact

OpenStudy (anonymous):

so basically when x= -1/11 or you can say whenever you multiply 11 by -1/11 and then add 2 you would get 1 as your final answer

OpenStudy (anonymous):

i think u r ryt

OpenStudy (anonymous):

bt any1 have another ansr

OpenStudy (anonymous):

i usually work backwards and use opposite functions because its easier

OpenStudy (anonymous):

at leats when dealing with algebra

OpenStudy (anonymous):

frnds hows this ADD 2 HOURS TO 11 O'CLOCK AND YOU GET 1 O'CLOCK

OpenStudy (anonymous):

that works but are we talking time or are you doing algebra

OpenStudy (anonymous):

well ash its just a puzzle

OpenStudy (anonymous):

wait i thought it was 11 and 2 not -10 we were talking abo ut

OpenStudy (anonymous):

oh lord im talking to a math geek that gets kick doing math while im struggling through my work sheet

OpenStudy (anonymous):

srry *kicks

OpenStudy (anonymous):

@KRAZZy there are a lot of answers your question isnt worded right i believe

OpenStudy (anonymous):

pls tell me just 1 ansr

OpenStudy (anonymous):

thats how the math sheet is take a look exercise 1

OpenStudy (anonymous):

like you could also say 11+2 = 13 and then divide it by 13 to get 1 but just did what i did before

OpenStudy (anonymous):

bt @hamza_b23 u hv any solution without put x :P

OpenStudy (anonymous):

yeah whenever 11 is mulitplied by -1/11 and then you add 2 you get 1

OpenStudy (anonymous):

ok dude thnx for help @hamza_b23